# Aritro Pathak

### v1

@@ 7 Future directions
If we consider the setting of [Tol24], where the boundary $\partial\Omega$ is contained in a co-dimension one hyperplane $\Gamma$, then one considers corkscrew balls contained within this hyperplane and the nearest ‘corner’ point $q_{1}$ on $\partial\Omega\cap\Gamma$ to this corkscrew point $p_{1}$ in question. There it was shown that the dimension drop phenomenon works for boundaries of dimension at least $n-3/2-\epsilon$ for some small enough epsilon. In future course of study, one expects to use the methods of this paper to study this question.
@@ 8 Acknowledgements
The author is grateful to Stephen Montgomery-Smith and Ridhhipratim Basu for helpful discussions.

### v2

@@ 7 Future directions
If we consider the setting of [Tol24], where the boundary $\partial\Omega$ is contained in a co-dimension one hyperplane $\Gamma$, then one considers corkscrew balls contained within this hyperplane and the nearest ‘corner’ point $q_{1}$ on $\partial\Omega\cap\Gamma$ to this corkscrew point $p_{1}$ in question. There it was shown that the dimension drop phenomenon works for boundaries of dimension at least $n-3/2-\epsilon$ for some small enough epsilon. In future course of study, we hope to use the methods of this paper to study this question.
@@ 8 Acknowledgements
The author is grateful to Stephen Montgomery-Smith and Ridhhipratim Basu for helpful discussions. The figure was generated using Chatgpt 5.